OpenStudy (anonymous):

Please help Find two consecutive odd numbers such that the sum of one-fifth of the smaller and four-sevenths of the larger is equal to fifty-nine. A.)73 and 75 B.)75 and 77 C.)77 and 79

OpenStudy (softballgirl372015):

59=1/59x + 4/7(x+2) x=first number x+2= consecutive odd number

OpenStudy (anonymous):

okay thanks. @Softballgirl372015

OpenStudy (softballgirl372015):

np :) happy to help!